OpenStudy (anonymous):

can anybody help me in week 10 hw and lab 10.plz

OpenStudy (anonymous):

chek ur msgs

OpenStudy (anonymous):

H10P1 r*u q*v p*w s*t

OpenStudy (anonymous):

H10P2 ((j*w*C)*(R1+R2) - (w*w*L*C)+1)/(j*w*C) inf inf (L*j*w)/((1-L*C*w^2)+(L/(R1+R2))*j*w) 0 0 ((R1+(1/(j*w*C)))*(R2+j*w*L))/((R1+(1/(j*w*C)))+(R2+j*w*L)) R2 R1

OpenStudy (anonymous):

H10P3 R2/((R1+R2-L*C*R1*w^2) + (R1*R2*C + L)*j*w) -(R2 + L*j*w)/((L*C*w^2 - 1) - R2*C*j*w) (1/(R1*w))*sqrt((R1-R2)/R2) (1/w)*sqrt(R1*R2-R2^2) 1057.01 52.85

OpenStudy (anonymous):

if last two answers doesn't work use this Cmatch=(sqrt (R1/R2 -1)) / (w * R1) Lmatch=((sqrt (R1/R2 -1)) * R2) / w

OpenStudy (anonymous):

LAB 10 ???

OpenStudy (anonymous):

Lab 10 1) 1/(R*C) 2) R/L 3) L 4)800e-6 5) C 6) 20e-9 For the last answer i have no idea how that works XD !

OpenStudy (anonymous):

For last question we can use a RLC circuit with f1 = 1KHz and f2=1MHz L=R/(f1*2*pi) C=1/(f2*R*2*pi) I would like to know how do you find: 4)800e-6 6) 20e-9 I tried with the same formula above but maybe i cannot find the right frequency!!!

OpenStudy (anonymous):

Hey this is Lab 10 solution

OpenStudy (anonymous):

WEEK 11??